Construct the confidence interval for the population standard deviation for the given values. Round your answers to one decimal place. n=23, s=9.9, and c=0.95

Could someone help teach me the formula to solve this in Excel! Thanks so much!
Answer #1

alpha = 1 - 0.95 = 0.05

The critical values for alpha = 0.05 and df = 22 are chi^2(1-alpha/2,n-1) = 10.9823 and chi^2(alpha/2,n-1) = 36.7807

value for upper limit = CHISQ.INV(0.05/2,22) = 10.9823
value for lower limit = CHISQ.INV.RT(0.05/2,22) = 36.7807

lower endpoint = sqrt(22*9.9^2/36.7807) = 7.7

upper endpoint = sqrt(22*9.9^2/10.9823) = 14.0